Assume that a procedure yield a binomial distribution with n =6 trials and a probability of success of p =0.90. Use a binomial probability table to find the probability that the number of successes × is exactly 2

omial Probabilties Table Binomal Probablitos Print Done

uiz 6 of 10 (4corrpkite) ▼ Binonial Probabilities Table Print Done ii..@E@하하하하하하-6
Answer #1

n = Number of trials = 6

p = Probability of success in a single trial = 0.90

So,

q = 1 - p =0.10

Table of the Binomial Cumulative Distribution gives:

For p = 0.90, n = 6, x = 2; Cumulative Probability from X =0 to X =2, we get : 0.001

For p =0.90,n = 6, x = 1, Cumulative Probability from X =0 to X = 1: we get: 0.000

Thus,

The probability that the number of successes X is exactly 2 = P(X = 2) is given by:
P(X = 2) = 0.001 - 0.000 = 0.001

So,

Answer is:

0.001